i have some physics questions.

If an infant can scream at a sound level of 111.0-dB measured at a distance 11.0-cm from their mouth, how far away should a parent stand so that the child can be seen but not heard? (The threshold of hearing is 0-dB, or equivalently I0 = 10-12-Wm-2.) Assume spherical symmetry for the dispersion of the sound wave energy (this is of course a bad assumption).

Answer km

How much power is the child putting in to screaming?

Answer

Answer=__________ W
Answer #1

We know that, decibel intensity is given by

Db = 10 log(I/Io)

111 = 10 log(I/10^-12)

11.1 = log I + 12

0.9 = log I

I = 10^0.9 = 7.94

7.94*0.11^2/R^2 = 10^-12

R = sqrt(7.94e12 *0.11^2)

= 310000 m

= 310 km

B) power = 7.94* 4 pi*0.11^2

= 1.207 W